osama1k:

Martin needs 3.25 gallons of red paint for his small barn and 1.25 gallons of paint for his shed. He spends $81 on paint. How much does the paint cost per gallon?

Moonrosegold:

3.25 + 1.25 = 4.5 $81 / 4.5 = $18.00 4.3 + 1.7 = 6 $41.94 / 6 = $6.99

jhonyy9:

@Moonrosegold please do you can explain your above wrote work ? thank you in advance

Moonrosegold:

Okay so if your wondering how much the paint costs per gallon for Martin, then you add 3.25 to 1.25 and it equals 4.5. Then you divide it with $81 and it will equal $18.00. then you add 4.3 to 1.7 which equals 6. And last but not least you divide $41.94 to 6 and you will get the total of $6.99 as your total cost.

Moonrosegold:

And also your welcome.

jhonyy9:

pardon more exactly the 81 divide by 4,5 yes ?

Moonrosegold:

Yes.

jhonyy9:

sorry but where you see in the text of posted problem this details with ,, then you add 4.3 to 1.7 which equals 6. And last but not least you divide $41.94 to 6 and you will get the total of $6.99 as your total cost." i dont see them nowhere